OpenStudy (anonymous):

factor. 2p^3 + 6p^2 + 3p + 9

OpenStudy (mathstudent55):

Factor by grouping. Factor out a common factor out of the first two terms. Factor out a common factor out of the last two terms. 'What do you get?

OpenStudy (anonymous):

Solve each quadratic equation. Show your work. 14. (2x – 1)(x + 7) = 0 15. x + 3x = 10

OpenStudy (anonymous):

i solved it on my own. can you help with these though?

OpenStudy (mathstudent55):

In 14. the polynomial is already factored and set equal to zero. Just set each factor equal to zero and solve.

OpenStudy (anonymous):

im confused

OpenStudy (mathstudent55):

14. (2x – 1)(x + 7) = 0 2x - 1 = 0 or x + 7 = 0 Solve each equation above.

OpenStudy (anonymous):

what

OpenStudy (mathstudent55):

In 14., a product can only equal zero if one of its factors is zero. You are given the product of two factors (2x - 1) and (x + 7) is equal to zero.

OpenStudy (mathstudent55):

The only way a product can equal zero is if at least one of the factors is equal to zero.

OpenStudy (anonymous):

so thats the answer?

OpenStudy (mathstudent55):

(2x – 1)(x + 7) = 0 Set 2x - 1 = 0 and set x + 7 = 0 Now solve each equation above.

OpenStudy (mathstudent55):

Can you solve 2x - 1 = 0 ?

OpenStudy (anonymous):

add 1 to both sides and divide by 2?

OpenStudy (mathstudent55):

Yes. Then solve x + 7 = 0

OpenStudy (anonymous):

-2?

OpenStudy (mathstudent55):

The solution is the two answers you got separated by the word "or".

OpenStudy (anonymous):

-7

OpenStudy (mathstudent55):

2x - 1 = 0 2x = 1 x = 1/2 x + 7 = 0 x = -7 Solution: x = 1/2 or x = -7

OpenStudy (anonymous):

thankyou. number 15?

OpenStudy (mathstudent55):

In 15. is the first x supposed to be x^2, or is it correct the way you have it?

OpenStudy (mathstudent55):

You wrote: 15. x + 3x = 10

OpenStudy (anonymous):

its x^2+3x=10

OpenStudy (mathstudent55):

Ok. The first step is to subtract 10 from both sides to have all terms on the left side equaling zero.

OpenStudy (anonymous):

but the 10 is separate already

OpenStudy (mathstudent55):

Yes, but you don't want it separate. You need it on the left side as -10. You need just zero on the right side.

OpenStudy (anonymous):

x^2+3x-10=0??

OpenStudy (mathstudent55):